Sunday, April 24, 2011

Nightlog 2011-04-23

SA: Alexei
SO: Siphelo
Others: Eric, Charl, Peter, Serafim


=> We have got 22 SALTICAM images of TPyx tonight.
I do not understand how we were able to get these images
taking into account all problems around.

=> The telescope software looks VERY UNSTABLE current days.
It looks like general communication problem.

Many things for SCAM interface did not work tonight,
but maybe because general communication problem.

We have problems with payload from the evening and were need
to initialize it. There were problems but finally it was done
with help of SALT team.
During observations of TPyx, we have tried to move RSS guide
probe, that results payload major fault. SALT team was unable
to recover this problem.

Finally, Siphelo was able to start payload MMI, but SALTICAM
filters stopped to work and gripper timout...

=> So, I think that currently we need software people here
to have telescope efficiency above -0.

=> Brilliant weather outside. Seeing is about 1".


No comments:

Post a Comment